OpenStudy (anonymous):

What is the mean of the following data values? 22, 37, 49, 15, 92 PLEASE I NEED HELP!!!

OpenStudy (anonymous):

mean is the sum of all the value of the datas which is divided by the number of data 22+37+49+15+92/5 =215/5= 43

OpenStudy (triciaal):

the mean is the average find the sum of all the values and divide by the number of terms that you had to add